Sam wants to fence his rectangular garden plot along its length on the front side. The perimeter of the plot is 24 meters, and its length is 3 times its width.
If x is the length and y is the width of the plot, set the equations of the system.
Solve and determine the values of x and y.
If the cost of fencing is $25 per meter, how much does Sam need to spend on fencing?

Answer by ankor@dixie-net.com(22740)

Sam wants to fence his rectangular garden plot along its length on the front side.
The perimeter of the plot is 24 meters, and its length is 3 times its width.
x = 3y
If x is the length and y is the width of the plot, set the equations of the system.
2x + 2y = 24
simplify, divide by 2
x + y = 12
Solve and determine the values of x and y.
replace x with 3y
3y + y = 12
4y = 12
y = 3 meters is the width
then
3(3) = 9 meters is the length
:
If the cost of fencing is $25 per meter, how much does Sam need to spend on fencing?
Assuming he wants of only fence the front length
9 * 27 = $243